OpenStudy (anonymous):

Find the area under the curve y=3-x^2 and above the x-axis. Round to three decimal places. So would I start off finding the antiderivative right? if so, that would be 3x - (x^3/3) ? @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

good, start \[\Large \int(3 - x^2)dx = 3x - \frac{1}{3}x^3 + C\]

jimthompson5910 (jim_thompson5910):

it doesn't give you the endpoints, odd

jimthompson5910 (jim_thompson5910):

oh nvm

jimthompson5910 (jim_thompson5910):

you need to find the x-intercepts

OpenStudy (anonymous):

ohhhh okay haha is this one different than the last one? :/

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):